OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). 2 sin2x = sin x

OpenStudy (anonymous):

Just need to see what the correct answer is, my answer choices are: A. x = pi/3 , 2pi/3 B. x = pi/2 , 3pi/2 , pi/3 , 2pi/3 C. x = 0 , pi , pi/6 , 5pi/6 D. x = pi/6 , 5pi/6